OpenStudy (keana):

Which point lies in the solution set for the following system of inequalities? x + y > 2 2x - y < 6 a: (4, 1) b: (3, -1) c: (0, 0) d: (1, 3)

OpenStudy (helder_edwin):

just replace the points u r given and see what happens. for instance: if (x,y)=(3,-1) then x=3 and y=-1 then \[ \large x+y=3+(-1)=2>2 \] is this correct?